Hook Industries's capital structure consists solely of debt and common equity. It can issue debt at rd = 10%, and its common stock currently pays a $2.50 dividend per share (D0 = $2.50). The stock's price is currently $21.50, its dividend is expected to grow at a constant rate of 7% per year, its tax rate is 35%, and its WACC is 12.10%. What percentage of the company's capital structure consists of debt? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Cost of equity=(D1/Current price)+Growth rate

=[(2.5*1.07)/21.5]+0.07

=19.4418605%

Cost of debt after-tax=10*(1-tax rate)

=10*(1-0.35)=6.5%

Let debt be $x

Equity be $y

Total=(x+y)

WACC=Respective cost*Respective weight

12.1=(x/(x+y)*6.5)+(y/(x+y)*19.4418605)

12.1*(x+y)=6.5x+19.4418605y

12.1x+12.1y=6.5x+19.4418605y

x=(19.4418605-12.1)y/(12.1-6.5)

=1.31104652 y

Total=x+y

=2.31104652 y

Hence weight of debt=1.31104652/2.31104652

=56.73%(Approx)